Suppose the marginal propensity to consume is 0.8.

The government increases government spending

and taxes by $10 billion. What happens to

aggregate output demanded?
Answer #1

MPC = 0.8

Government spending increases by 10 billion so G = 10

Y/G = 1/(1- MPC)  

Y/10 = 1/(1 - 0.8)

Y/10   = 1/0.2

Y = 5(10)

= 50

NOW  T = 10

YT = - MPC(1 - MPC)

= - 0.8/(1 - 0.8)

= - 0.8/0.2

Y/10 = - 4

Y = -40

overall effect = 50 + (-40)

= 10  

which means aggregate output increases by 10 billion